OpenStudy (anonymous):

Write the first trigonometric function in terms of the second for θ in the given quadrant. sec θ, tan θ, θ in quadrant II

Nnesha (nnesha):

don't forget this identity \[\huge\rm \sin^2\theta +\cos^2 \theta=1\]

Nnesha (nnesha):

sec x =1/cosx tanx = sinx/cosx cot= cosx /sin x csc x= 1/sinx sec ,csc cot AND 1/sin ,1/cos,1/ tan are reciprocal of each other

Nnesha (nnesha):

to find sec theta first solve that identity for cos theta

Nnesha (nnesha):

\[\huge\rm \sin^2\theta +\cos^2 \theta=1\] cos theta = ?

Nnesha (nnesha):

 remember the CAST rule for signs